MONT 108N – Mathematics Through Time Oral Presentations on Euclid’s Elements October 29, 2010

General Information

As announced in the course syllabus, as a part of our reading of Book I of Euclid’s Elements, you will be making an oral presentation to the class (together with a classmate, or two classmates in one case – 17 is “unfortunately prime(!)”) on one of the Propositions.

1) You should plan to use about 15 minutes for your part of this, leaving a couple of minutes for questions at the end, if possible.

2) Your presentation should explain the statement of the Proposition and its proof in detail in your own words – don’t just copy the text of the Proposition and its proof. In particular give the dependences on previously proved propositions to justify each step of the proof. It’s OK to use modernized language and concepts like congruence that Euclid doesn’t really have in our form.

3) If there is some particularly interesting aspect of the Proposition, or if the statement is a common fact that you know from high school geometry, or if it relates closely to some of the other Propositions, you will want to mention that as well. (The notes on David Joyce’s web site are very good for this sort of thing!)

4) Each person in the team should give a part of the presentation – part of the idea of Montserrat is to start developing your public speaking abilities in addition to your writing.

5) You can give the presentation either as a “chalk talk” or with PowerPoint if you prefer. If you choose PowerPoint, though, please make it really basic(!) No fancy slide transitions or audio effects, please! (I can help out with the mechanics of displaying a PowerPoint presentation in Swords 209 if you need that.)

6) See me during office hours if you need help deciphering the statement or proof, or if you would like to run a presentation by me before you do “the real thing” in class.

7) I will adjust the due date of the second paper so that it comes after this week. Watch for an announcement about this.

Schedule, Assigned Teams and Propositions

1) Monday, November 8:

a) Lucia and Sarah – Proposition I.33

b) Catherine and Dan – Proposition I.34

c) Muhammad and Kyle – Proposition I.35

2) Wednesday, November 10:

a) Matt and Astrid – Proposition I.36

b) John and Andrew – Proposition I.37

c) Xuan, Kaitlin, Max – Proposition I.38 3) Friday, November 12:

a) Tom and Kylee – Proposition I.41 b) Kevin and Peter – Proposition I.46